OpenStudy (anonymous):

If x>1, then (1-x)/(1-sqrt. x)= A. x B. sqrt. x C. x- sqrt. x D. 1+sqrt. x E 1- sqrt. x

OpenStudy (anonymous):

\[(1-x)/(1-\sqrt{x})=\]

OpenStudy (anonymous):

\[\large \frac{1-x}{1-\sqrt{x}}=\frac{1-x}{1-\sqrt{x}}\times \frac{1+\sqrt{x}}{1+\sqrt{x}}\] \[\large =\frac{(1-x)(1+\sqrt{x})}{1-x}\] \[\large =1+\sqrt{x}\]

OpenStudy (anonymous):

why did you multiple it by 1+ sqrt. x/ 1+ sqrt. x ????

OpenStudy (anonymous):

That's called rationalising the denominator.

OpenStudy (anonymous):

You usually do that to get rid of square root signs. You multiply the denominator with it's conjugate.

OpenStudy (anonymous):

but sign is different?

OpenStudy (anonymous):

I already explained it to you in the second sentence?

OpenStudy (anonymous):

nooo, i mean it is +, cause it is not 1+ sqrt. x, but 1- sqrt. x?

OpenStudy (anonymous):

You multiply by its conjugate in this scenario because it only has one term under a square root sign. It's not all under a square root sign. For example: \[\large \frac{1+x}{\sqrt{1+x}}\] You can rationalise the denominator by multiplying the denominator with the same thing. But for questions like the above, you use the property of "difference between two squares". For example: \[\large (x-1)(x+1)=x^2-1\] And if it was square root x: \[\large (\sqrt{x}-1)(\sqrt{x}+1)=(\sqrt{x})(\sqrt{x})-(1)(1)\] \[\large =x-1\]

OpenStudy (anonymous):

The conjugate of \[1+\sqrt{x}\] is \[1-\sqrt{x}\]

OpenStudy (anonymous):

When you multiply by it's conjuagte, you get rid of the square root sign instantly.

OpenStudy (anonymous):

OK, understand. thank you very much :)))